OpenStudy (anonymous):

How can I Solve this ?? Find the values of x for which f(x)=g(x) 1.) f(x)= x^2 + 3x + 4 , G(x)= 5x + 28 PLS. HELP ME :))))))) and show the solution :)

OpenStudy (anonymous):

I'm assuming g(x) and G(x) are the same thing here? If so, then x^2+3x+4=5x+28 x^2-2x-24=0 (x-6)(x+4)=0 x=6, x=-4 You can now plug them back in and check.

OpenStudy (anonymous):

Thank You :D

OpenStudy (anonymous):

How come that x +4 became x= -4 ?? and x-6 became x=6 ??